(5.1.15) Refer to the accompanying​ table, which describes results from groups of 8 births from 8 different sets of parents. The random variable x represents the number of girls among 8 children. Find the mean and standard deviation for the number of girls in 8 births.
Number of
Girls x ​P(x)
0 0.006
1 0.035
2 0.109
3 0.222
4 0.277
5 0.213
6 0.107
7 0.026
8 0.005
The mean is mu ? ​girl(s)
1
Expert's answer
2020-11-25T19:25:33-0500

The following table shows the provided outputs of the discrete random variables, along with the corresponding probabilities:



Therefore, the population mean is calculated as follows:

"\mu =\sum_{i=1}^{n} x_{i}*P(x_i)"


=0*0.006+1*0.035+2*0.109+3*0.222+4*0.277+5*0.213+6*0.107+7*.026+8*0.005


=3.956



Standard deviation:

We first compute the expected value of "X^2" .

"E(X^2)=\sum_{i=1}^{n} x_{i}^2*P(x_i^2)"

="0 \n2\n *0.006+1 \n2\n *0.035+2 \n2\n *0.109+3 \n2\n *0.222+4 \n2\n *0.277+5 \n2\n *0.213+6 \n2\n *0.107+7 \n2\n *.026+8 \n2\n *0.005=\n17.672\n\u200b"


Therefore, the population variance is computed as follows:

"\u03c3 ^\n2\n \n\u200b\t\n \n=\n\n\u200b\t\n \nE(X \n^2\n )\u2212E(X) ^\n2"

= "17.672\u22123.956 ^\n2" = 2.0221


And finally, taking square root to the variance we get that the population standard deviation is "\sigma = \sqrt{ \sigma^2} = \sqrt{2.0221} = 1.422"